Show your calculations for the followings;

1) energy of a photon of frequency of 5.20x1011 s-1.
Answer #1

E = hV

h = 6.625*10^(-34) j.s V = 5.20*1011 s-1.

E = 6.625*10^(-34)*5.20*10^11 = 3.445*10^-22 joule.